# Building Information Modelling MRR modular reconfigurable robot

### v1

@@ V-D Discussion
A compactness-robustness tradeoff can clearly be seen in both mean and Pareto-optimal scores. The results shown in Figure 3b indicate a similar tradeoff concerning the reconfiguration time. When starting from an initially assembled robot from a previous task, the first feasible robot for drilling was identified within three minutes. However, if high robustness is essential, a reconfiguration time of over 20 minutes is necessary to assemble the corresponding robot [^7] Note that in general, setting up a new modular robot takes around half that time. In this experiment, the most robust solutions required a complete disassembly of an existing configuration before, resulting in a long reconfiguration time.. Both results underline the benefits of multi-objective optimization: Based on the expected disturbances, the required compactness, and the time at hand, an operator can choose the situationally optimal solution for deployment.
@@
The robot could adjust for mobile base positioning errors in all real-world experiments. As expected, the adjustment increases with higher positioning errors. Further, as shown in Figure 5, the absolute errors differ between runs, and for five out of six experiments, the necessary adjustment was below 50% of the maximum adjustment. If a more compact (and thereby less robust) robot along the Pareto-optimal designs needs to be chosen, high positioning errors of the base could likely be overcome by reapproaching the wall and repeating the calibration.
@@
The primary purpose of the conducted experiments was to demonstrate the feasibility of the proposed framework. The gains of the controller were intentionally set to a lower level to prevent excessive force, resulting in reduced precision. This ensures that the robot does not push too far into the wall, even if the vision system fails to accurately detect the marker. Still, as shown in Figure 6, we observed a relative horizontal precision of $1$ cm. In future applications, the system could be enhanced with a more accurate vision system to improve its accuracy. Additionally, implementing a force control mechanism would further robustify the system and enable a more precise and adaptive interaction with the environment.
VI Conclusion
@@
We developed and integrated a holistic framework allowing construction workers to instruct a mobile modular robot arm for building construction. In contrast to existing solutions, the modularity of hardware and software components and multi-objective optimization of robot morphology enables tailoring the proposed approach to specific needs in situ. By integrating BIM information with optimization and control frameworks from the robotics domain, we have created a holistic framework that requires minimal intervention and robotics expertise for deployment. Further, our approach is not limited to geometric data of the used BIM models; future work could incorporate information about material properties to enable automation even for dynamically more challenging tasks, such as impact drilling.

### v2

@@ V-E Discussion
When starting from an assembled robot from a previous task, the mechanical assembly of the first valid robot takes three minutes. However, if high robustness is essential, a reconfiguration time of more than 20 minutes is necessary to assemble the corresponding robot [^6] Note that in general, setting up a new modular robot takes around half that time. In this experiment, the most robust solutions required complete disassembly of an existing configuration, resulting in a long reconfiguration time.. Both results underline the benefits of multiobjective optimization: Based on the expected disturbances, the required compactness, and the time at hand, an operator can choose the situationally optimal solution for deployment.
The results of our simulations show the clear benefit of algorithmic support in this process, as some of the outcomes contradict initial human intuition. When adapting a robot to be more robust against calibration errors, the first reflex would often consist of adding more joints. As indicated by the observed trade-off between compactness and robustness, a correlation exists between more modules and higher robustness. However, our experiments have shown that the additional mass and lever of a new module often lead to a violation of the joint torque constraint. Further, the changed kinematics usually resulted in a complete recomputation of the motion plan, which is often less robust due to self-collisions. Our results show that the use of long passive links often leads to high robustness, as they increase the reach of the robot while adding only a fraction of the mass of a new joint. Thus, to achieve a more robust robot, a complete reconfiguration might be necessary, as further highlighted in Figure 3b. Design patterns, such as a minimum required number of joints or an almost consistent choice for the first two joints, emerged, indicating the existence of specific dominant configurations. However, as demonstrated by the spray painting simulation, these patterns are task-specific.
@@
The robot could adjust for mobile base positioning errors in all real-world experiments. As expected, the adjustment increases with higher positioning errors. Furthermore, as shown in Figure 5, the absolute errors differ between runs, and for five of the six experiments, the necessary adjustment was below 50% of the maximum adjustment. If a more compact (and thereby less robust) robot along the Pareto front needs to be chosen, high positioning errors of the base could likely be overcome by reapproaching the wall and repeating the calibration.
@@
The primary purpose of the experiments conducted was to demonstrate the feasibility of the proposed framework. The control gains were intentionally set to a lower level to prevent excessive force, resulting in reduced precision. This ensures that the robot does not push too far into the wall, even if the vision system fails to accurately detect the marker. Still, as shown in Figure 6, we observed a relative horizontal precision of $1$ cm. In future applications, the system could be enhanced with a more accurate vision system to improve its accuracy. Furthermore, implementing a force control mechanism would further robustify the system and enable a more precise and adaptive interaction with the environment.
VI Conclusion and Limitations
@@
We developed and integrated a holistic framework that allows construction workers to instruct a mobile modular robot arm for building construction. In contrast to existing solutions, the modularity of hardware and software components and multiobjective optimization of robot morphology enables tailoring the proposed approach to specific needs in situ. By integrating BIM information into optimization and control frameworks from the robotics domain, we have created a holistic framework that requires minimal intervention and robotics expertise for deployment. In addition, our approach is not limited to considering geometric data of the used BIM models; future work could also incorporate information about material properties to enable automation even for dynamically more challenging tasks, such as impact drilling.
Evolutionary approaches do not provide guarantees regarding their optimality or completeness. In addition, the quality of the solutions depends not only on our method but also on the underlying motion planner used for trajectory generation. Lastly, a successful transfer to the real world depends on the quality of the BIM data and the sensors. While the overall approach applies to various tasks, future applications, such as bricklaying or plastering, will require adjusted simulation models, constraints, and optimization objectives.